# LoFi: Scalable Local Image Reconstruction with Implicit Neural Representation

### v1

@@ III LoFi: a local field processor
We then consider a local forward operator ${\mathbf{A}}$ where the intensity ${\mathbf{f}}[x,y]$ at pixel $(x,y)\in\mathbb{R}^{2}$ relates only to a small neighborhood of $(x,y)$ in ${\mathbf{q}}$. Based on this locality hypothesis, we design a coordinate-based neural network inspired by implicit neural representation.
To recover the image at the target pixel $(x,y)$ individually, we extract local information around $(x,y)$ from the observed image ${\mathbf{q}}$ containing $K$ pixels. We denote the (flattened) extracted information by ${\mathbf{p}}_{x,y}\in\mathbb{R}^{K\times C}$. As illustrated in Figure 1, we process ${\mathbf{p}}_{x,y}$ using a neural network $\text{NN}_{\theta}:\mathbb{R}^{K\times C}\to\mathbb{R}^{C}$ with parameters $\theta$ to approximate the image intensity ${\mathbf{f}}[x,y]$,
@@
This coordinate-based image representation enables image recovery at any continuous coordinate with small memory. We call our framework LoFi (Local Field). As we typically use a small neighborhood size $K$, we can parametrize $\text{NN}_{\theta}$ using an MLP. It is worth noting that a similar framework is used in  (Khorashadizadeh et al., 2023a) for continuous image super-resolution using CNNs. In the following sections, we provide further details regarding the LoFi architecture and introduce the pre-processing filter, patch geometry, and training strategy.
@@ III-A MultiMLPs for large neighborhoods
One important factor in LoFi architecture is the neighborhood dimension $K$; larger $K$ can provide more detailed information to improve reconstructions. However, by increasing $K$, the size of MLP quadratically scales resulting in a bulky network and slow training. To address this issue, we propose MultiMLP, an alternative architecture to the current MLP to handle large neighborhoods. Inspired by vision transformers (Dosovitskiy et al., 2020), we split the extracted neighboring pixels (here over a circular geometry) into smaller chunks, each processed by a separate MLP as shown in Figure 2. Their outputs are then mixed by another MLP.
@@ III-B1 Differentiable patch extraction
While we can use the standard square patch to extract the local information, it does not provide a differentiable operator from the input coordinate to the pixel intensity. To enable image reconstruction at arbitrary continuous coordinates, we extract the local information ${\mathbf{p}}_{x,y}$ using bicubic interpolation, providing a continuous differentiable operator.
@@ III-B2 Learnable patch geometry
The differentiable patch extraction suggested above also brings important flexibility; it allows us to learn the patch geometry centered around the target coordinate $(x,y)$ by first defining a set of learnable coordinate offsets,
@@
We emphasize that the coordinates $I_{(x,y)}$ need not be aligned with the pixel grid as we use differentiable bicubic interpolation to enable off-the-grid evaluation. Unlike most other work using square patches, we initialize $\Delta I$ to a circular geometry; this results in better rotation invariance as discussed in Section S-II-B in the supplementary materials.
@@ III-C Coordinate-conditioned patch geometry
While the learnable patch geometry introduced in (9) can provide additional flexibility to the network, the learned patch geometry is fixed for all the pixels in the image. In practice, the location of the relevant information changes for different pixels. To enable coordinate-conditioned patch geometry (CCPG), we use another neural network (typically a simple MLP) $\text{CCPG}_{\psi}:\mathbb{R}^{K\times C}\to\mathbb{R}^{2K}$ which takes a local patch around pixel $(x,y)$ and estimates the position of the coordinates inside the patch as follows,
@@
We then evaluate ${\mathbf{q}}$ at $(x,y)+\Delta I_{CCPG}(x,y)$. The learned patch geometry $\Delta I_{CCPG}(x,y)$ can localize the position of the relevant information for image recovery at pixel $(x,y)$ for corrupted image ${\mathbf{q}}$. It is worth noting that a similar strategy has been developed by (Dai et al., 2017) for object detection using CNNs.
@@
where in each iteration $1\leq i\leq T$, we feed noisy image ${\mathbf{q}}$ evaluated at the estimated patch position in the previous step $p_{x,y}^{(i-1)}$ to a separate neural network $\text{CCPG}_{\psi}^{(i)}$. As we will show in our experiments in Section V-B, the learned patch geometry indeed improves over iterations.
@@ III-D Noise suppression filter
The observed image ${\mathbf{q}}$ is often corrupted with significant noise making the inversion challenging for the $\text{NN}_{\theta}$ in (8). To address this, we apply pre-processing filters ${\mathbf{h}}\in\mathbb{R}^{s\times s\times L}$ to the observed image before patch extraction,
@@
where $\star$ denotes the convolution operator, $s$ is the filter size and $L$ is the number of filters. This process is reminiscent of filtering techniques commonly used in computed tomography imaging (Diwakar and Kumar, 2018). We can also concatenate the filtered and noisy images along the channel dimension.
@@
where $\mathbf{F}$ and $\mathbf{F}^{-1}$ are the forward and inverse 2D Fourier transforms, and the filter ${\mathbf{H}}$ is treated as a learnable parameter derived from the data. The filter ${\mathbf{H}}$ can have large support in image space, capturing global information, in contrast to the local convolutional operator in (14). This large filter support can allow the $\text{NN}_{\theta}$ to overfit global features and compromise model generalization. To mitigate this issue, we initialize the filter ${\mathbf{H}}$ with an all-one Fourier representation, which corresponds to the smallest support in pixel space. During training, the filter expands its support based on the image resolution, noise level, and data distribution while keeping the filter support small. To further control support size, we could add a regularization penalty based on the filter support but in subsequent experiments presented in Section S-II-C in the supplementary materials, we noticed that the learned filter has small support even without any regularization.
@@ III-E Resolution-agnostic memory usage in training
At inference time, we can recover the image at any target pixel ${\mathbf{x}}$ from the observed image ${\mathbf{q}}_{\text{test}}$ as $\hat{{\mathbf{f}}}_{\text{test}}({\mathbf{x}})=\text{LoFi}_{\phi^{\star}}({\mathbf{x}},{\mathbf{q}}_{\text{test}})$. From (16), it is clear that LoFi can be trained on mini-batches of both objects and pixels, leading to nearly constant memory usage during training and inference. This flexibility also allows us to train LoFi on datasets with images in various resolutions.
IV Further Applications
Figure 3: Performance comparison on LDCT (30dB noise) in resolution $128\times 128$.
@@ IV-A Learning the imaging geometry
Many seemingly complex image-to-image transforms, like those involving translations and rotations, alter image content using simpler coordinate transformations. For example, transposing a 2D image causes significant pixel differences in high-dimensional space; but this transformation relies on a simple, low-dimensional coordinate swap. When CNNs are fitted on image-to-image tasks, they are directly trained to model the changes in pixel values and, therefore, tend to overlook the coordinate transformations that underlie the pixel value changes. This is particularly problematic for large-scale changes in the coordinate system as the CNN should have a large receptive field to capture these geometric changes. This limitation has been explored by existing works to build deep neural networks that can model the changes in the coordinate system as well as the image content. For example, the authors of (Kothari et al., 2020) separately model changes of the coordinate system and the image content using CNNs.
@@
We train the above architecture end-to-end where the task of INR is to learn the changes in the coordinate system, while LoFi learns the changes in the image content from the corrected coordinate system. We call this new architecture INR-LoFi and compare its performance with CNNs for a toy experiment in Section V-C.
@@ IV-B Low-resolution priors for high-resolution unsupervised image reconstruction
As shown in Section III-E, the proposed LoFi is a supervised learning framework that provides fast image recovery from observations. However, supervised learning methods necessitate paired training data comprising measurements and corresponding target images which is hard to obtain in many real scientific problems (Jin et al., 2017; McCann et al., 2017; Liu et al., 2022; Wei and Chen, 2018; Khorashadizadeh et al., 2023c). Additionally, the related models require re-training even for small changes in the forward model.
To address this issue, unsupervised methods learn a data-driven prior from target images. Deep generative models have been extensively used for solving imaging inverse problems where a pre-trained image generator is used as the prior (Bora et al., 2017; Kothari et al., 2021; Asim et al., 2020; Kawar et al., 2022; Khorashadizadeh et al., 2023b; Liu et al., 2023). However, training deep generative models often requires large-scale datasets and substantial computational resources (Kingma and Dhariwal, 2018; Ho et al., 2020).
An alternative is to use a Gaussian denoiser as a prior in a plug-and-play (PnP) framework (Venkatakrishnan et al., 2013; Chan et al., 2016; Romano et al., 2017). PnP algorithms achieved SOTA performance across a variety of image reconstruction tasks including image super-resolution and deblurring (Dong et al., 2018; Zhang et al., 2021), computed tomography (CT) (Wei et al., 2022) and magnetic resonance imaging (MRI) (Ahmad et al., 2020; Wei et al., 2020); see also the recent review (Kamilov et al., 2023). Since PnP algorithms decouple the likelihood from the image prior they enable the use of pre-trained denoisers. Compared to generative priors they are often cheaper in both training and inference. Recently, the authors of (Zhu et al., 2023) leveraged diffusion models (Ho et al., 2020; Dhariwal and Nichol, 2021) as a pre-trained denoiser in PnP framework and achieved excellent performance for various image restoration tasks. We provide a brief overview of the PnP-ADMM framework in Section S-I-E of the supplementary material.
In this section, we use a pre-trained LoFi denoiser as a prior in the plug-and-play (PnP) framework to benefit from its continuous image representation for solving inverse problems; this integration allows us to use a LoFi denoiser trained on low-resolution images for solving inverse problems at any higher resolution.
We employ a pre-trained LoFi denoiser in PnP-ADMM and call this framework LoFi-ADMM described in Algorithm 1. Remarkably, while we train LoFi denoiser on low-resolution images, LoFi-ADMM can conveniently run at any arbitrary higher resolutions thanks to the continuous image representation of LoFi.
@@ V Experiments
Figure 4: Comparative analysis for image denoising in resolution $512\times 512$ where different models are trained on a tiny dataset with 9 training samples. The PSNR of the reconstructed test samples is demonstrated per iterations during training. CNNs in particular multiscale versions show severe overfitting while LoFi shows a robust convergence and significantly outperforms CNNs thanks to its locality design.
Figure 5: Performance comparison on image denoising where different models are trained for 20000 iterations on a tiny training set with 9 images shown in Figure S1.
We run extensive experiments to evaluate the performance of LoFi from different aspects. In Section V-A, we compare LoFi with successful CNNs for several image reconstruction tasks. In Section V-B, we use our proposed method in Section III-C for analysis and interpretation of the reconstructed image. In Section V-C, we showcase the success of INR-LoFi and the failure of CNNs on a simple toy experiment where we have changes in the coordinate system. We compare LoFi with CNNs in terms of memory usage and training time in Section V-D. In Section V-E, we apply our unsupervised framework LoFi-ADMM to two inverse problems, image in-painting and radio interferometry in cosmology.
@@ V-A Scalable image reconstruction with LoFi
We assess the performance of LoFi on inverse problems with local forward operators including low-dose computed tomography (LDCT), image denoising, and dark matter mapping in cosmology. As baselines, we consider standard CNNs including two single-scale models DnCNN (Zhang et al., 2017a) and IRCNN (Zhang et al., 2017b) and two multiscale models, U-Net (Ronneberger et al., 2015) and DRU-Net (Zhang et al., 2021). We assess the reconstruction quality with peak signal-to-noise ratio (PSNR) and Structural Similarity Index (SSIM) (Wang et al., 2004). Further details regarding the LoFi architecture and training details are available in Section S-I in the supplementary materials.
Low-dose computed tomography:
We simulate parallel-beam low-dose CT with $r=180$ projections uniformly distributed around the object. The measured projections are contaminated with 30dB Gaussian noise. We use 1000 training samples of chest images from the LoDoPaB-CT dataset (Leuschner et al., 2021) in resolution $128\times 128$. The model performance is assessed on 100 in-distribution test samples of chest images and we use 16 brain samples (Hssayeni et al., 2020) to evaluate model generalization on out-of-distribution (OOD) data.
The upper row of Figure 3 illustrates the performance of different models on in-distribution chest images. Interestingly, LoFi can achieve comparable performance with standard CNNs by using only MLP modules. The lower row demonstrates the model performance on out-of-distribution (OOD) brain samples. LoFi significantly outperforms CNNs, indicating the strong generalization achieved by the locality of the architecture.
TABLE I: Model comparison on the reconstruction quality for LDCT (30dB SNR) averaged on 100 in-distribution chest and 16 OOD brain samples.
@@
DnCNN (Zhang et al., 2017a) 34.2 0.90 27.6 0.82
IRCNN (Zhang et al., 2017b) 34.5 0.91 29.8 0.93
U-Net (Ronneberger et al., 2015) 33.0 0.90 13.9 0.81
DRU-Net (Zhang et al., 2021) 33.8 0.90 29.0 0.90
LoFi (ours) 33.9 0.90 30.6 0.94
Image denoising:
@@
In this experiment, we consider image denoising where we use a very small dataset of only 10 images at a resolution $512\times 512$ —9 images for training and 1 for testing, shown in Figure S1 in the supplementary materials. The images are normalized to a range between 0 and 1, and zero-mean Gaussian noise with $\sigma=0.15$ is added. We compare LoFi with various CNNs with a comparable number of weights (3M). Figure 4 presents the PSNR of the reconstructed test image across training iterations for each model. While strong CNNs such as DRU-Net, U-Net, and DnCNN initially perform well, they quickly overfit the training data, leading to a sharp drop in test image performance. IRCNN demonstrates improved generalization but also begins to overfit after 10000 iterations. Remarkably, LoFi shows strong generalization due to its locality design, avoiding overfitting even on this small dataset. This experiment clearly shows that we can conveniently train LoFi on a tiny dataset without being concerned about overfitting and needing early stopping or other regularization techniques. Final reconstructions after 20000 iterations are illustrated in Figure 5.
Dark matter mapping:
As shown in (7), the dark matter mapping problem can be interpreted as a denoising problem characterized by a local forward operator. We apply LoFi on the naive KS inversion $\kappa_{\text{KS}}$ to recover the convergence field $\kappa$. For additional details regarding the experimental setup, please refer to Section S-I-C in the supplementary materials.
@@
In Figure 6, the reconstructed convergence fields from the KS image are presented for various models at resolution $128\times 128$. The performance parity between LoFi and CNNs indicates the validity of the locality property derived from (7).
@@
Figure 6: Performance comparison on reconstructed convergence fields for dark matter mapping in resolution $128\times 128$.
V-B Reconstruction analysis and interpretation

### v2

@@ III LoFi: a local field processor
We then consider a local forward operator ${\mathbf{A}}$ where the intensity ${\mathbf{f}}[x,y]$ at pixel $(x,y)\in\mathbb{R}^{2}$ relates only to a small neighborhood of $(x,y)$ in ${\mathbf{q}}$. Based on this locality hypothesis, we design a coordinate-based neural network inspired by implicit neural representations.
To recover the image (only) at the target pixel $(x,y)$, we extract a neighborhood of $K$ pixels around $(x,y)$ from the observed image ${\mathbf{q}}$. We denote the (flattened) extracted pixels by ${\mathbf{p}}_{x,y}\in\mathbb{R}^{K\times C}$. As illustrated in Figure 2, we process ${\mathbf{p}}_{x,y}$ using a neural network $\text{NN}_{\theta}:\mathbb{R}^{K\times C}\to\mathbb{R}^{C}$ with parameters $\theta$ to approximate the image intensity ${\mathbf{f}}[x,y]$,
@@
This coordinate-based image representation enables image recovery at any continuous coordinate with small memory. We call our framework LoFi (Local Field). As we typically use a small neighborhood size $K$, we can parametrize $\text{NN}_{\theta}$ using an MLP. In the following sections, we provide further details regarding the LoFi architecture and introduce the pre-processing filter, patch geometry, and training strategy.
@@ III-A MultiMLPs for large neighborhoods
One important factor in LoFi architecture is the neighborhood dimension $K$; larger $K$ can provide more detailed information to improve reconstructions. However, the size of the MLP quadratically scales with $K$ resulting in large networks and slow training. To address this issue, we propose MultiMLP, an alternative architecture to handle large neighborhoods. Inspired by vision transformers [50], we split the extracted neighboring pixels (here over a circular geometry) into smaller chunks, each processed by a separate MLP as shown in Figure 3. Their outputs are then mixed by another MLP.
@@ III-B1 Differentiable patch extraction
While we could use the standard square patch to extract the local information, this would not result in a differentiable map from the input coordinate to the pixel intensity. To enable image reconstruction at arbitrary continuous coordinate, we extract ${\mathbf{p}}_{x,y}$ using bicubic interpolation which results in a differentiable architecture.
@@ III-B2 Learnable patch geometry
This differentiable patch extraction opens up the possibility to learn the patch geometry centered around the target coordinate $(x,y)$ by first defining a set of learnable coordinate offsets,
@@
We emphasize that the coordinates $I_{(x,y)}$ need not be aligned with the pixel grid as we use differentiable bicubic interpolation to enable off-the-grid evaluation. Unlike most other work using square patches, we initialize $\Delta I$ to a circular geometry; this results in better rotation invariance as discussed in Section VIII-B in the supplementary materials.
@@ III-C Coordinate-conditioned patch geometry
The learned patch geometry in (9) is fixed. In practice, the relative locations of the most relevant information changes for different pixels. To enable coordinate-conditioned patch geometry (CCPG), we use another neural network (typically a simple MLP) $\text{CCPG}_{\psi}:\mathbb{R}^{K\times C}\to\mathbb{R}^{2K}$ which takes a local patch around pixel $(x,y)$ and estimates the position of the coordinates inside the patch as follows,
@@
We then evaluate ${\mathbf{q}}$ at $(x,y)+\Delta I_{CCPG}(x,y)$. The learned patch geometry $\Delta I_{CCPG}(x,y)$ can localize the position of the relevant information for image recovery at pixel $(x,y)$ for corrupted image ${\mathbf{q}}$. A similar strategy has been developed by [68] for object detection using CNNs.
@@
where in each iteration $1\leq i\leq T$, we feed noisy image ${\mathbf{q}}$ evaluated at the estimated patch position in the previous step $p_{x,y}^{(i-1)}$ to a separate neural network $\text{CCPG}_{\psi}^{(i)}$. As we will show in our experiments in Section V-C, the learned patch geometry indeed improves over iterations.
@@ III-D Noise suppression filter
The observed image ${\mathbf{q}}$ is corrupted with noise. Optimal filters for correlated noise can have a large receptive field which cannot be implemented by the localized $\text{NN}_{\theta}$ in (8). To address this, we apply pre-processing convolutional filters ${\mathbf{h}}\in\mathbb{R}^{s\times s\times L}$ to the observed image before patch extraction,
@@
where $\star$ denotes the convolution operator, $s$ is the filter size and $L$ is the number of filters. This process is reminiscent of filtering techniques commonly used in computed tomography imaging [69]. We can also concatenate the filtered and noisy images along the channel dimension.
@@
where $\mathbf{F}$ and $\mathbf{F}^{-1}$ are the forward and inverse 2D Fourier transforms, and the diagonal matrix ${\mathbf{H}}$ is a learnable parameter fitted to the data. The resulting filter can have large support in image space, capturing global information. We initialize the filter ${\mathbf{H}}$ with all ones in the Fourier domain (an identity filter) which corresponds to the smallest support in pixel space. During training, the filter expands its support based on the image resolution, noise level, and data distribution.
@@ III-E Resolution-agnostic memory usage in training
At inference time, we can recover the image at any target pixel ${\mathbf{x}}$ from the observed image ${\mathbf{q}}_{\text{test}}$ as $\hat{{\mathbf{f}}}_{\text{test}}({\mathbf{x}})=\text{LoFi}_{\phi^{\star}}({\mathbf{x}},{\mathbf{q}}_{\text{test}})$. From (16), it is clear that LoFi can be trained on mini-batches of both objects and pixels, enabling an almost constant-in-resolution memory usage during training and inference. This flexibility also allows us to train LoFi on datasets with images in various resolutions.
IV Further Extensions and Applications
Figure 4: Performance comparison on LDCT (30dB noise) at resolution $128\times 128$ for in-distribution (chest) and OOD (brain) data.
@@ IV-A Learning the imaging geometry
Many seemingly complex image-to-image transforms, like those involving translations and rotations, alter image content using simpler coordinate transformations. For example, transposing a 2D image causes significant pixel differences in the high-dimensional pixel space; but the transformation relies on a simple, low-dimensional coordinate swap. When CNNs are fitted on image-to-image tasks, they are directly trained to model the changes in pixel values and, therefore, tend to overlook the coordinate transformations that underlie the pixel value changes. This is particularly problematic for large-scale changes in the coordinate system as the CNN should have a large receptive field to capture these geometric changes. This limitation has been explored by existing works to build deep neural networks that can model the changes in the coordinate system as well as the image content. For example, the authors of [70] separately model changes of the coordinate system and the image content using CNNs.
@@
We train the above architecture end-to-end where the task of INR is to learn the changes in the coordinate system, while LoFi learns the changes in the image content from the corrected coordinate system. We call this new architecture INR-LoFi and compare its performance with CNNs for a toy experiment in Section V-D.
@@ IV-B Low-resolution priors for high-resolution unsupervised image reconstruction
As shown in Section III-E, the proposed LoFi is a supervised learning framework and requires paired training data of measurements and corresponding target images. This is hard to obtain in many real scientific problems [6, 71, 72, 73]. Additionally, supervised models require re-training even for small changes in the forward model.
To address this issue, unsupervised methods learn a data-driven prior from only target images. Once trained, they can be used as a prior for solving any inverse problem using an iterative process. Unsupervised methods based on deep generative models [74, 32, 75, 76, 77] and plug-and-play (PnP) [78, 79, 80] show strong performance on various image reconstruction tasks. We provide a brief overview of the PnP-ADMM framework in Section VII-E of the supplementary material.
In this section, we use a pre-trained LoFi denoiser as a prior in the plug-and-play (PnP) framework. This integration allows us to use a LoFi denoiser trained on low-resolution images to solve inverse problems at any higher resolution. Since we use the PnP-ADMM backbone we call the resulting scheme LoFi-ADMM (cf. Algorithm 1). Even though the LoFi denoiser is trained on low-resolution images, LoFi-ADMM can be applied at arbitrary higher resolutions thanks to LoFi’s continuous image representation.
@@ V Experiments
We now present a suite of experimental results which illustrate the various aspects of LoFi. In Section V-A, we compare LoFi with successful CNNs and ViTs on several image reconstruction tasks. We compare LoFi with CNNs and ViTs in terms of memory usage and training time in Section V-B. In Section V-C, we use our proposed method in Section III-C for analysis and interpretation of the reconstructed image. In Section V-D, we showcase the success of INR-LoFi and the failure of CNNs on a simple toy experiment where we have changes in the coordinate system. In Section V-E, we apply our unsupervised framework LoFi-ADMM to two inverse problems, image in-painting and radio interferometry.
@@ V-A Scalable image reconstruction with LoFi
We assess the performance of LoFi on inverse problems with local operators including low-dose computed tomography (LDCT) upon applying FBP, image denoising, and dark matter mapping in cosmology. As baselines, we consider two standard single-scale CNNs, DnCNN [11] and IRCNN [45] and two multiscale CNNs, U-Net [5] and DRU-Net [15]. We also consider two strong vision transformers, Restormer [53] and SwinIR [12], which achieved state-of-the-art performance on various image reconstruction problems. We assess the reconstruction quality with peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM) [81]. Further details regarding the network architectures and training details are given in Section VII in the supplementary materials.
Low-dose computed tomography (LDCT)
We simulate parallel-beam LDCT with $r=180$ uniformly distributed projection orientations. We add white Gaussian noise to obtain the signal-to-noise (SNR) ratio of 30dB. We use 1000 training samples of chest images from the LoDoPaB-CT dataset [82] at resolution $128\times 128$. All models take FBP as input and return improved reconstruction in the output. The model performance is assessed on 64 in-distribution test samples of chest images. We additionally use 16 brain CT samples [83] to evaluate model generalization on out-of-distribution (OOD) data.
The upper row of Figure 4 illustrates the performance of different models on in-distribution chest images. We see that LoFi achieves performance comparable to strong CNNs and ViTs while using only simple MLP modules. The bottom row shows model performance on OOD brain samples. It is remarkable that LoFi significantly outperforms CNNs and ViTs, indicating the strong generalization achieved by the locality of the architecture. Quantitative results are presented in Table I.
TABLE I: Model comparison on the reconstruction quality for LDCT (30dB SNR) averaged on 64 in-distribution chest and 16 OOD brain samples.
@@
DnCNN [11] 34.2 0.90 27.6 0.82
IRCNN [45] 34.5 0.91 29.8 0.93
U-Net [5] 33.0 0.90 13.9 0.81
DRU-Net [15] 33.8 0.90 29.0 0.90
Restormer  [53] 33.5 0.89 28.9 0.79
SwinIR  [12] 34.5 0.90 29.3 0.94
LoFi (ours) 34.1 0.90 30.6 0.94
Image denoising with tiny datasets
Here we train LoFi on image denoising on a small dataset of only 10 images at resolution $512\times 512$; cf. Figure 13 in the supplementary materials. 9 images are used for training and 1 for evaluation. The images are scaled between 0 and 1 and we add white Gaussian noise with $\sigma=0.15$. We compare LoFi with CNN and Restormer models with a comparable number of trainable parameters (3M). SwinIR was excluded from this experiment due to its high computational cost and very long processing time at this resolution.
@@
Figure 5 presents the PSNR of the reconstructed test image across training iterations for each model. While CNNs initially perform well, they quickly overfit the training data, leading to a sharp drop in test image performance. Restormer and IRCNN demonstrate better generalization but also begin to overfit after 2500 and 10000 iterations, respectively. LoFi on the other hand generalizes strongly, avoiding overfitting even with 3M parameters and a tiny dataset. This is without any explicit regularization or early stopping. Final reconstructions after 20000 iterations are illustrated in Figure 6. Please also see Section VIII-A in the supplementary materials for additional experiments showcasing LoFi’s performance when trained on a large-scale dataset (CelebA-HQ) with high-resolution images.
Figure 5: Comparative analysis for image denoising at resolution $512\times 512$ where different models are trained on a tiny dataset with 9 training samples. The PSNR of the reconstructed test samples is demonstrated per iterations during training. CNNs, in particular multiscale versions, show severe overfitting while LoFi shows a robust convergence and significantly outperforms CNNs thanks to its locality design.
Figure 6: Performance comparison on image denoising where different models are trained for 20000 iterations on a tiny training set with 9 images shown in Figure 13.
Dark matter mapping
@@
We finally apply LoFi on the naive KS inversion $\kappa_{\text{KS}}$ to recover the convergence field $\kappa$. In Figure 7, the reconstructed convergence fields from the KS image are presented for various models at resolution $128\times 128$. The performance parity between LoFi and baselines, especially ViTs which can process long-range dependencies, indicates the validity of the locality property derived from (7). In Figure 17, we also demonstrate the learned filter ${\mathbf{H}}$ in (15); this figure can be useful to understand how it processes input KS images.
@@
Figure 7: Performance comparison on reconstructed convergence fields for dark matter mapping at resolution $128\times 128$.
V-B Computational efficiency
(a) Memory footprint
(b) Inference time
Figure 8: The memory and time requirements during inference for different models
As discussed in Section III-E, LoFi can be trained and run on machines with small GPUs by reducing the pixel batch size in both training and inference. Figure 1 compares memory usage and training time for 100 iterations with batch size 1 across different models and image resolutions. In this experiment, we used pixel batch size 512 for LoFi during training. This analysis shows that while the required memory and training time for CNNs and ViTs scales with image resolution, LoFi’s memory usage and training time remains nearly constant enabling efficient training on high-resolution images.
Moreover, LoFi training is fast even for large images. For example, training on $1024\times 1024$ images using 29900 samples from the CelebA-HQ dataset was completed in only 8 hours (200 epochs) on a single A100 GPU, requiring only 15 GB memory with object and pixel batch sizes 64 and 512 respectively. Additional details and qualitative results for this experiment can be found in Section VIII-A.
While LoFi is efficient during training, inference on large images can become expensive due to coordinate-based image synthesis. This requires iteratively reconstructing large images, as generating all pixels at once would need a bigger memory. Figure 8 compares the memory usage and inference time of different architectures. In this experiment, LoFi generates 16384 pixels per iteration. We mention potential strategies to improve inference efficiency in Section VI.
V-C Reconstruction analysis and interpretation